an airplane made a 2040-kilometer trip in the direction of the wind in 5 hours last tuesday. if the trip against the direction of the wind took an hour longer, find the rate of the plane in still air and the velocity of the wind.
Answer by stanbon(75887)

an airplane made a 2040-kilometer trip in the direction of the wind in 5 hours last tuesday. if the trip against the direction of the wind took an hour longer, find the rate of the plane in still air and the velocity of the wind.
--------------
Let "p" and "w" represent the plane speed and the wind speed.
--------------
With-wind DATA:
distance = 2040 km ; time = 5 hrs ; rate = 2040/5 = 408 kph
--------------
Against-wind DATA:
distance = 2040 km ; time = 6 hrs ; rate = 2040/6 = 340 kph
--------------
EQUATIONS:
p + w = 408
p - w = 340
-------------
Add to get:
2p = 748
p = 374 kph (speed of the plane in still air)
------------
Substitute into p+w = 408 to solve for w:
374 + w = 408
w = 34 kph (speed of the wind)
==================================
Cheers,
Stan H.